OpenStudy (anonymous):

A local hamburger shop sold a combined total of 642 hamburgers and cheeseburgers on Thursday. There were 58 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Thursday?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

263 cheeseburgers?

OpenStudy (johnweldon1993):

Not quite...because there were only 58 fewer cheeseburgers...lets try it again c + h = 642 c = h - 58 (h - 58) + h = 642 2h - 58 = 642 2h = 700 h = 350 So 350 hamburgers....and that means there would be 292 Cheeseburgers...since 350 - 58 = 292.. and to check....350 + 292 = 642...There we go :)

OpenStudy (anonymous):

Thanks so muchh

OpenStudy (johnweldon1993):

Of course! Anytime :)